# thanks: Corresponding author

### v2

@@ Abstract
Using the recently proposed covariant framework of general relativistic stochastic mechanics and stochastic thermodynamics, we proved the detailed and integral fluctuation theorems in curved spacetime. The time-reversal transformation is described as a transformation from the perspective of future-directed observer to that of the corresponding past-directed observer, which enables us to maintain general covariance throughout the construction. The result presented in this work may be applied in understanding the origin of irreversibility in macroscopic processes in the presence of relativistic gravity, as in most astrophysical or cosmological processes.
@@
Introduction. – A fundamental problem in modern statistical physics is the emergence of macroscopic irreversibility in systems which have time-reversal symmetry (TRS) in the underlying microscopic description. The relevant researches can be traced back to Boltzmann, who employed a scattering model with TRS in deriving the H-theorem [1], which states that the entropy of a macroscopic system cannot decrease in the course of time. However, the H-theorem has been under debates and challenges ever since its birth. The most famous challenge is known as Loschmidt paradox [2, 3], which argues that, if an entropy-increasing process exists for a system, the underlying TRS should also permit a corresponding entropy-decreasing process. The fluctuation theorems, proposed in the 1990s [4, 5, 6], offer an interpretation for Loschmidt paradox. Basically, these theorems state that, while both entropy-increasing and entropy-decreasing processes are permissible, the former is more probable.
@@
Relativistic Langevin dynamics. – Langevin equation describes the motion of a heavy particle, referred to as the Brownian particle, under the random disturbance of a heat reservoir. Conventionally, Langevin equation can be expressed in the form of Newton’s second law, incorporating elements such as the random force, damping force, and various other external forces. For a relativistic particle of mass $m$ and charge $q$ moving in $(d+1)$-dimensional spacetime $\mathcal{M}$ with the metric $g_{\mu\nu}$, the Langevin equation employing the particle’s proper time $\tau$ as evolution parameter is is referred to as LEτ [19],
@@
Conclusion. – Based on the covariant framework of general relativistic stochastic mechanics, we formulated the first law of stochastic thermodynamics in curved spacetime. By properly addressing the description for the TRS and its breaking, the corresponding fluctuation theorem is also proved in both the differential and integral forms. The observers and their behaviors under the TRT play a key role in our construction. The fact that the final fluctuation theorems (34) and (General Relativistic Fluctuation Theorems) take the same form as in non-relativistic context [8] is expected, because both the trajectory probability and the total entropy production are neither coordinate dependent nor observer dependent. The version of the fluctuation theorems proved in this work may shed some light on understanding the origin of irreversibility in the presence of relativistic gravitational interaction, as in most cases of astrophysical or cosmological processes.

### v3

@@ Abstract
Using the recently proposed covariant framework of general relativistic stochastic mechanics and stochastic thermodynamics, we proved the detailed and integral fluctuation theorems in curved spacetime. The time-reversal transformation is described as a transformation from the perspective of future-directed observer to that of the corresponding past-directed observer, which enables us to maintain general covariance throughout the construction.
@@
Introduction. – A fundamental problem in modern statistical physics is the emergence of macroscopic irreversibility in systems which have time-reversal symmetry (TRS) in the underlying microscopic description. The relevant researches can be traced back to Boltzmann, who employed a scattering model with TRS in deriving the H-theorem [1], which states that the entropy of a macroscopic system cannot decrease in the course of time. However, the H-theorem has been under debates and challenges ever since its birth. The most famous challenge is known as Loschmidt paradox [2, 3], which argues that, if an entropy-increasing process exists for a system, the underlying TRS should also permit a corresponding entropy-decreasing process. Nowadays, Loschmidt paradox is understood to be originated from the use of the molecular chaos hypothesis. After about one hundred and twenty years since the birth of the H-theorem, a number of fluctuation theorems were proposed [4, 5, 6], which provide an alternative quantitative description for the irreversibility of macroscopic systems by use of some equalities instead of the inequality presented by the H-theorem, and meanwhile attribute the origin of the irreversibility to the dissipative effects on the microscopic level.
@@
Relativistic Langevin dynamics. – Langevin equation describes the motion of a heavy particle, referred to as the Brownian particle, under the random disturbance of a heat reservoir. Conventionally, Langevin equation can be expressed in the form of Newton’s second law, incorporating elements such as the random force, damping force, and various other external forces. For a relativistic particle of mass $m$ and charge $q$ moving in $(d+1)$-dimensional spacetime $\mathcal{M}$ with the metric $g_{\mu\nu}$, the Langevin equation employing the particle’s proper time $\tau$ as evolution parameter is referred to as LEτ [19],
@@
Conclusion. – Based on the covariant framework of general relativistic stochastic mechanics, we formulated the first law of stochastic thermodynamics in curved spacetime. By properly addressing the description for the TRS and its breaking, the corresponding fluctuation theorem is also proved in both the differential and integral forms. The observers and their behaviors under the TRT play a key role in our construction. The fact that the final fluctuation theorems (34) and (General Relativistic Fluctuation Theorems) take the same form as in non-relativistic context [8] is expected, because both the trajectory probability and the total entropy production are neither coordinate dependent nor observer dependent.
The proof of the fluctuation theorems presented in this work extends the range of applicability of the fluctuation theorems to the cases involving strong gravity and heat reservoir with a non-uniform temperature. We expect that such extension may be found useful in certain astrophysical processes. We hope to come back on this point in future works.